# Order Out of Noise and Disorder: Fate of the Frustrated Manifold

### v1

@@ Abstract
[…]
@@ 1 Introduction
Unlike conventional spontaneous symmetry breaking, the symmetry-breaking direction is not frozen but slowly drifts via thermal noise, while the reduced-dimensional structure persists. On the sphere, this behavior constitutes a classical realization of dissipative Nambu-Goldstone modes, specifically the type-A diffusion modes in the Minami-Hidaka classification [2, 3], with diffusive dynamics ($\langle\Delta\theta^{2}\rangle\propto t$). The torus and cylinder, however, exhibit anomalous MSD scaling (superdiffusive with $\alpha\approx 1.7$ and ballistic with $\alpha\approx 2$, respectively), revealing that topology and boundary conditions can qualitatively modify collective dynamics.
@@ 2.4 Rotational Dynamics and Orientation Diffusion
where $\boldsymbol{\tau}_{\xi}(t)$ is the stochastic torque from thermal fluctuations. This is an Ornstein-Uhlenbeck process with relaxation time $\tau_{L}=I/\gamma$. As shown by van Kampen [19], in the overdamped limit ($\gamma\to\infty$ or equivalently $\tau_{L}\to 0$), the angular momentum becomes a “fast variable” that thermalizes instantaneously. The physically meaningful slow variables are the orientation angles, which undergo rotational diffusion. Appendix B provides a detailed derivation.
For our multi-particle system in the overdamped regime, we therefore focus on the collective orientation rather than angular momentum. On each manifold, the emergent structure (band, rings, or clusters) has an orientation that can be parameterized by angles. The natural observable is the cumulative rotation
@@
where $\omega(t)$ is the instantaneous angular velocity of the collective structure. This cumulative rotation undergoes a random walk when the deterministic torque vanishes.
Zero-Torque Property and Orientation Diffusion.
The zero-torque result (23) implies that pairwise interaction forces produce no deterministic drift in the collective orientation. In the language of rotational diffusion, the orientation evolves according to
@@
with no drift term. Here $D_{\text{rot}}$ is the effective rotational diffusion coefficient and $dW$ is a standard Wiener process. The mean-squared angular displacement therefore grows linearly:
@@
This diffusive scaling ($\alpha=1$ in $\langle\Delta\Theta^{2}\rangle\propto t^{\alpha}$) is expected when thermal torques are temporally uncorrelated. Deviations from $\alpha=1$ signal persistent correlations induced by the evolving particle configuration or global constraints from boundaries and metric structure.
@@
On all three manifolds, the collective orientation evolves with no deterministic drift from pairwise interactions. For uniform particle distributions with uncorrelated thermal noise, one expects diffusive scaling ($\alpha=1$). On the sphere, this is confirmed: the band orientation performs a random walk on the Goldstone manifold $S^{2}\cong\text{SO}(3)/\text{SO}(2)$, with ensemble-averaged MSD yielding $\alpha\approx 0.9$. However, the cylinder and torus exhibit anomalous scaling (ballistic with $\alpha\approx 2$ and superdiffusive with $\alpha\approx 1.7$, respectively), indicating that the non-uniform particle distribution after structure formation induces persistent correlations in the effective thermal torques.
@@ 3.6 Results on the Sphere: $S^{2}\to S^{1}$
The band geometry emerges from energy minimization in the frustrated interaction landscape, not from angular momentum dynamics. Particles concentrate along a great circle because this configuration minimizes the total potential energy for typical realizations of the random couplings. The zero-torque property (23) ensures that the band orientation, once established, evolves only through thermal noise with no deterministic drift. The cumulative rotation $\Theta(t)$ tracks this stochastic orientation evolution.
@@
Quantitative Rotational Dynamics Analysis.
Figure 2 presents an analysis of rotational dynamics and structure formation on the sphere for a simulation with $N=400$ particles at temperature $T=0.4$, covering the full time range from $t=0$ to $t=50$.
Figure 2: Sphere $S^{2}$: Rotational dynamics analysis for $N=400$ particles, $T=0.4$, Gaussian $\phi_{ij}$, full time series $t=0$ to $50$. (a) Total cumulative rotation $\Theta_{z}(t)$ showing large-amplitude fluctuations; vertical dashed line marks band formation at $t\approx 5$. (b) Comparison of total $\Theta_{z}$ (blue) and band contribution (green), demonstrating correlation after formation. (c) $\Theta_{z}$ distribution in steady state with $\langle\Theta_{z}\rangle=-9.6$, $\sigma=148.7$. (d) Angular velocity $\omega_{z}$ distribution with $\sigma=0.37$. (e) Rotational energy $E_{\text{rot}}=\Theta_{z}^{2}/(2I)$ time series. (f) Formation phase detail (t=0t=01515). (g) $\Theta_{z}$ autocorrelation with rapid decay ($\tau<0.2$). (h) Band vs. total $\Theta_{z}$ scatter showing correlation (slope $\approx 0.42$, $r\approx 0.57$). (i) Angular MSD with single-run exponent $\alpha\approx 1.4$; ensemble average over 5 runs yields $\alpha=0.9\pm 0.3$ (diffusive). (j) Phase comparison: formation phase (red, $\sigma=205$) vs. steady state (blue, $\sigma=149$).
Cumulative Rotation Statistics.
The cumulative rotation $\Theta_{z}(t)=\int_{0}^{t}\sum_{i}\rho_{i}^{2}\,d\theta_{i}$ (defined in Eq. (28)) exhibits large-amplitude fluctuations characteristic of a random walk. Table 4 summarizes the statistics for both the formation and steady-state phases.
Table 4: Cumulative rotation statistics on the sphere separated by phase (dimensionless units). The formation phase shows larger fluctuations as particles reorganize; the steady state exhibits reduced variance once the band has formed.
@@
Full simulation t=0t=05050 9.4-9.4 $155.4$
The mean cumulative rotation is small compared to its fluctuations ($|\langle\Theta_{z}\rangle|\ll\sigma_{\Theta_{z}}$), consistent with the rotational symmetry of the system: there is no preferred direction for net rotation when averaged over initial conditions, though any single realization may have a nonzero mean. The large fluctuations reflect the stochastic nature of Langevin dynamics, where thermal noise continuously drives orientation changes. The theoretical prediction (67) gives $d\Theta_{z}=\sqrt{2D}\sum_{i}\sin\theta_{i}\,dW_{i}^{\phi}$, implying that $\Theta_{z}$ undergoes pure diffusion with no deterministic drift. For uniformly distributed particles, the effective diffusion coefficient is $D_{\text{rot}}=2ND/3$ (Table 1), predicting $\sigma_{\Theta_{z}}\sim\sqrt{2D_{\text{rot}}\Delta t}\sim\sqrt{(4ND/3)\Delta t}$. With $N=400$, $D=0.4$, and effective observation time $\Delta t\sim 45$, this gives $\sigma\sim\sqrt{(4\times 400\times 0.4/3)\times 45}\approx 98$, which is of the correct order of magnitude given that the band configuration concentrates particles near the equator (where $\sin\theta\approx 1$), enhancing fluctuations beyond the uniform-distribution prediction. Notably, the standard deviation decreases from the formation phase ($\sigma=205$) to the steady state ($\sigma=149$), indicating that band formation partially constrains the rotational fluctuations.
Diffusive Band Rotation.
The mean-square displacement (MSD) of the cumulative azimuthal rotation $\theta(t)$ (Figure 2i) exhibits power-law scaling:
$\langle\Delta\theta^{2}\rangle\sim t^{\alpha},\quad\alpha\approx 1.1$ (36)
The exponent $\alpha\approx 1$ indicates approximately diffusive behavior, consistent with normal Brownian motion of the band orientation. This diffusive behavior is a direct consequence of the theoretical result (67): since $d\Theta_{z}=\sqrt{2D}\sum_{i}\sin\theta_{i}\,dW_{i}^{\phi}$ with no deterministic drift term, the cumulative rotation undergoes a pure random walk. The thermostat continuously exchanges rotational impulses with the system, leading to random-walk behavior of the cumulative rotation angle with $\langle\Delta\theta^{2}\rangle\propto D_{\text{rot}}t$, where $D_{\text{rot}}$ is the effective rotational diffusion coefficient from Table 1.
Quenched Disorder and Ensemble Averaging.
An important feature of the MSD analysis concerns the variation of the fitted exponent $\alpha$ across independent simulation runs. In any finite realization, the time-averaged angular velocity $\bar{\omega}=T^{-1}\int_{0}^{T}\omega(t)\,dt$ may be non-zero, contributing an apparent ballistic term $\bar{\omega}^{2}\tau^{2}$ to the MSD. This non-zero drift reflects genuine physics: each realization explores a different free energy landscape determined by the specific quenched interaction matrix $\phi_{ij}$.
The MSD of cumulative rotation decomposes as:
$\langle(\Delta\theta)^{2}\rangle=\mathrm{Var}(\Delta\theta)+[\mathrm{Mean}(\Delta\theta)]^{2}=\mathrm{Var}(\Delta\theta)+\bar{\omega}^{2}\tau^{2}$ (37)
where the variance term reflects diffusive spreading and the mean-squared term captures systematic drift. Physically, $\bar{\omega}$ encodes information about the instanton path the system followed during structure formation: different realizations of $\phi_{ij}$ create different potential landscapes with different barrier heights, saddle points, and most-probable transition paths.
We verified this interpretation by analyzing five independent simulations for each geometry, each with different random seeds generating different $\phi_{ij}$ matrices and initial conditions. On the sphere, individual runs showed $\alpha$ ranging from $0.55$ to $1.13$, with sample means $\bar{\omega}$ ranging from 0.034-0.034 to +0.010+0.010; some landscapes favor clockwise drift, others counterclockwise, with no preferred direction on average (ensemble mean consistent with zero, $p=0.14$). Ensemble averaging yielded $\alpha=0.888$, confirming that the sphere exhibits true diffusive dynamics ($\alpha\approx 1$), consistent with the theoretical prediction of pure rotational diffusion with no deterministic drift.
In contrast, the cylinder and torus exhibit qualitatively different behavior that persists under ensemble averaging, as detailed in Sections 3.7 and 3.8.
Autocorrelation Structure.
The $\Theta_{z}$ autocorrelation function (Figure 2g) shows rapid decay on a timescale $\tau<0.2$ time units, corresponding to the correlation time of velocity fluctuations. This rapid decorrelation indicates that successive cumulative rotation values become statistically independent on very short timescales, consistent with the white-noise character of Langevin dynamics at temperatures where the damping timescale is comparable to the thermal fluctuation timescale.
Band Contribution to Rotation.
In the steady state, essentially all particles ($\gtrsim 90\%$) are concentrated within the band structure visible in Figure 1(d). The correlation between band and total cumulative rotation (Figure 2h, slope $\approx 0.42$, $r\approx 0.57$) confirms that the organized band structure contributes significantly to the collective rotational dynamics. The slope being less than unity reflects the fact that particles at different latitudes within the finite-width band contribute differently to $\Theta_{z}$: particles near the band center (equator of the instantaneous rotation) contribute maximally, while particles at the band edges contribute less due to the $\sin\theta$ weighting.
Comparison with Glassy Relaxation.
The relaxation dynamics on the sphere exhibits features characteristic of glassy systems:
  • Two-stage relaxation: Fast initial dispersion (t=0t=022) followed by band assembly (t=2t=255).
  • Metastability: The system explores the dispersed configuration before finding the great-circle attractor.
  • Fluctuation reduction: The cumulative rotation standard deviation decreases from $\sigma=205$ (formation) to $\sigma=149$ (steady state), indicating increased dynamical constraints after band formation.

### v2

@@ Abstract
We study Langevin dynamics of $N$ Brownian particles on two-dimensional Riemannian manifolds, interacting through pairwise potentials linear in geodesic distance with quenched random couplings. These frustrated Brownian particles experience competing demands of random attractive and repulsive interactions while confined to curved surfaces. We consider three geometries: the sphere $S^{2}$, torus $T^{2}$, and bounded cylinder. Our central finding is disorder-induced dimension reduction with spontaneous rotational symmetry breaking: order emerges from two sources of randomness (thermal noise and quenched disorder), with manifold topology determining the character of emerging structures. Glassy relaxation drives particles from 2D distributions to quasi-1D structures: bands on $S^{2}$, rings on $T^{2}$, and localized clusters on the cylinder. Unlike conventional symmetry breaking, the symmetry-breaking direction is not frozen but evolves slowly via thermal noise. On the sphere, the structure normal precesses diffusively on the Goldstone manifold with correlation time $\tau_{c}\approx 18$, a classical realization of type-A dissipative Nambu-Goldstone dynamics. The model requires no thermodynamic gradients, no fine-tuning, and no slow external input. We discuss connections to spin glass theory, quantum field theory, astrophysical structure formation, and self-organizing systems. The model admits a large-$N$ limit yielding statistical field theory on Riemannian surfaces, while remaining experimentally realizable in colloidal and soft matter systems.
@@ 1 Introduction
Unlike conventional spontaneous symmetry breaking, the symmetry-breaking direction is not frozen but slowly drifts via thermal noise, while the reduced-dimensional structure persists. On the sphere, the structure normal $\hat{\mathbf{n}}(t)$ (the direction perpendicular to the band plane) slowly precesses on the Goldstone manifold $S^{2}\cong\text{SO}(3)/\text{SO}(2)$, analogous to the magnetization direction in a ferromagnet. This slow precession, with correlation time $\tau_{c}\approx 18$ far exceeding the microscopic thermal time, constitutes a classical realization of dissipative Nambu-Goldstone modes, specifically the type-A diffusion modes in the Minami-Hidaka classification [5, 6]. The torus and cylinder show more complex behavior due to their different symmetry breaking patterns and geometric constraints.
@@ 2.4 Rotational Dynamics and Orientation Diffusion
where $\boldsymbol{\tau}_{\xi}(t)$ is the stochastic torque from thermal fluctuations. This is an Ornstein-Uhlenbeck process with relaxation time $\tau_{L}=I/\gamma$. As shown by van Kampen [22], in the overdamped limit ($\gamma\to\infty$ or equivalently $\tau_{L}\to 0$), the angular momentum becomes a “fast variable” that thermalizes instantaneously. The physically meaningful slow variable is the orientation of the rigid body in 3D space, which undergoes slow rotational diffusion. Appendix B provides a detailed derivation.
Two Observables: Fast Cumulative Rotation and Slow Orientation.
For our multi-particle system, we analyze two distinct observables that capture different aspects of collective rotational dynamics:
  1. 1.
    Cumulative rotation $\Theta(t)$: The total amount of rotation accumulated around a fixed axis (the $z$-axis in our geometries). Since $\Theta(t)$ is driven by the instantaneous angular momentum, which is a fast variable in the overdamped limit, we expect $\Theta(t)$ to exhibit fast diffusive behavior similar to angular momentum itself. It grows without bound and does not represent a slow collective mode.
  2. 2.
    Structure normal $\hat{\mathbf{n}}(t)$: A unit vector specifying the instantaneous orientation of the emergent structure (band, rings, or clusters) in 3D space. This is the direct implementation of van Kampen’s “orientation angle” for our collective structures. Unlike $\Theta(t)$, the structure normal is a bounded quantity (it lives on $S^{2}$) and represents the true slow variable that undergoes rotational diffusion on a much longer timescale than the thermal velocity correlations.
While $\Theta(t)$ answers “how much has the structure rotated?”, $\hat{\mathbf{n}}(t)$ answers “where is the structure pointing now?” We analyze both observables in Section 3.4 to obtain a complete characterization of the rotational dynamics.
Cumulative Rotation: Definition and Expected Behavior.
We define the cumulative rotation as
@@
where $\omega(t)$ is the instantaneous angular velocity of the collective structure. The zero-torque result (23) implies that pairwise interaction forces produce no deterministic drift in $\Theta(t)$. In the absence of drift, the cumulative rotation evolves according to
@@
where $D_{\text{rot}}$ is the effective rotational diffusion coefficient and $dW$ is a standard Wiener process. The mean-squared displacement therefore grows linearly:
@@
This diffusive scaling ($\alpha=1$ in $\langle\Delta\Theta^{2}\rangle\propto t^{\alpha}$) is the expected “fast” behavior when thermal torques are temporally uncorrelated. Deviations from $\alpha=1$ signal persistent correlations induced by the evolving particle configuration or global constraints from boundaries and metric structure.
@@
On all three manifolds, the cumulative rotation $\Theta(t)$ evolves with no deterministic drift from pairwise interactions. For uniform particle distributions with uncorrelated thermal noise, one expects diffusive MSD scaling ($\alpha=1$) for this fast variable. On the sphere, this is confirmed: the cumulative rotation MSD yields ensemble-averaged $\alpha\approx 0.9$. However, the cylinder and torus exhibit anomalous scaling (ballistic with $\alpha\approx 2$ and superdiffusive with $\alpha\approx 1.7$, respectively), indicating that the non-uniform particle distribution after structure formation induces persistent correlations in the effective thermal torques. The structure normal $\hat{\mathbf{n}}(t)$ (the slow Goldstone mode variable) exhibits different dynamics characterized by the correlation time $\tau_{c}$, as described in Section 2.5.
2.5 Structure Orientation from Position Covariance
We now provide the technical definition of the structure normal $\hat{\mathbf{n}}(t)$ introduced above. This observable implements van Kampen’s slow orientation variable for our collective structures.
Definition via Eigendecomposition.
For each geometry, we define the structure normal $\hat{\mathbf{n}}(t)$ through the position covariance matrix in the 3D embedding space,
$M_{ij}=\frac{1}{N}\sum_{k=1}^{N}r_{k}^{(i)}r_{k}^{(j)},\quad i,j\in\{x,y,z\},$ (30)
where $\mathbf{r}_{k}=(r_{k}^{(x)},r_{k}^{(y)},r_{k}^{(z)})$ are the 3D Cartesian coordinates of particle $k$. For the sphere, these are simply the unit sphere coordinates; for the cylinder and torus, we use the standard embeddings $\mathbf{r}(\theta,z)=(R\cos\theta,R\sin\theta,z)$ and $\mathbf{r}(\theta,\varphi)=((R+r\cos\varphi)\cos\theta,(R+r\cos\varphi)\sin\theta,r\sin\varphi)$, respectively.
The eigendecomposition of $M$ yields three eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$ with corresponding eigenvectors. The structure normal $\hat{\mathbf{n}}$ is defined as the eigenvector associated with the smallest eigenvalue $\lambda_{3}$: this is the direction along which the particle distribution has minimal extent, i.e., the direction perpendicular to the plane of the band, ring, or cluster configuration. We define the structure thickness as
$\mathcal{T}=\sqrt{\frac{\lambda_{3}}{\lambda_{1}}},$ (31)
which measures the aspect ratio of the particle distribution. For a perfect planar configuration, $\mathcal{T}=0$; for an isotropic distribution, $\mathcal{T}=1$.
Expected Dynamics.
For a well-formed structure ($\mathcal{T}\ll 1$), the normal $\hat{\mathbf{n}}(t)$ lives on the unit sphere $S^{2}$. Its dynamics reflect the slow reorientation of the collective structure under thermal fluctuations. The autocorrelation function $C(\tau)=\langle\hat{\mathbf{n}}(t)\cdot\hat{\mathbf{n}}(t+\tau)\rangle$ defines a correlation time $\tau_{c}$ that measures how quickly the structure loses memory of its orientation. We expect $\tau_{c}$ to be much longer than the microscopic thermal correlation time, reflecting the adiabatic nature of orientation dynamics in the overdamped regime.
2.6 Summary: Two Complementary Observables
The theoretical framework developed in this section provides two complementary observables for characterizing collective rotational dynamics:
  1. 1.
    Cumulative rotation $\Theta(t)$: A fast, unbounded quantity tracking total angular displacement around a fixed axis. Driven by the instantaneous angular momentum, its mean-squared displacement $\langle\Delta\Theta^{2}\rangle\propto t^{\alpha}$ is expected to show diffusive scaling ($\alpha=1$) when thermal torques are uncorrelated.
  2. 2.
    Structure normal $\hat{\mathbf{n}}(t)$: The slow, bounded unit vector specifying instantaneous orientation in 3D. Its autocorrelation defines a correlation time $\tau_{c}\gg\tau_{\text{thermal}}$, and its power spectrum $S_{n_{z}}(\omega)$ probes the frequency content of orientation fluctuations.
In Section 3.4, we present numerical results for both observables across all three geometries. For each manifold, we first describe the structure formation process and then analyze the cumulative rotation and orientation dynamics in the non-equilibrium steady state.
@@ 3.6 Results on the Sphere: $S^{2}\to S^{1}$
The band geometry emerges from energy minimization in the frustrated interaction landscape, not from angular momentum dynamics. Particles concentrate along a great circle because this configuration minimizes the total potential energy for typical realizations of the random couplings. The zero-torque property (23) ensures that the band orientation, once established, evolves only through thermal noise with no deterministic drift. Two observables characterize this evolution: the cumulative rotation $\Theta_{z}(t)$ measures total angular displacement around the $z$-axis (a fast variable), while the structure normal $\hat{\mathbf{n}}(t)$ specifies the instantaneous band orientation in 3D (the slow Goldstone mode).
@@
Cumulative Rotation Dynamics.
The cumulative rotation $\Theta_{z}(t)=\int_{0}^{t}\sum_{i}\rho_{i}^{2}\,d\theta_{i}$ (defined in Eq. (28)) exhibits large-amplitude fluctuations characteristic of a random walk, as predicted by the theoretical result (65). Table 4 summarizes the statistics.
Table 4: Cumulative rotation statistics on the sphere separated by phase (dimensionless units). The formation phase shows larger fluctuations; the steady state exhibits reduced variance once the band has formed.
@@
The mean cumulative rotation is small compared to its fluctuations ($|\langle\Theta_{z}\rangle|\ll\sigma_{\Theta_{z}}$), consistent with rotational symmetry: there is no preferred direction for net rotation. The theoretical prediction $d\Theta_{z}=\sqrt{2D}\sum_{i}\sin\theta_{i}\,dW_{i}^{\phi}$ implies pure diffusion with no deterministic drift. The standard deviation decreases from the formation phase ($\sigma=205$) to the steady state ($\sigma=149$), indicating that band formation partially constrains the rotational fluctuations.
MSD Scaling and Ensemble Averaging.
The mean-square displacement of the cumulative rotation exhibits power-law scaling $\langle\Delta\theta^{2}\rangle\sim t^{\alpha}$. Ensemble averaging over five independent simulations (each with different quenched disorder $\phi_{ij}$) yields $\alpha=0.9\pm 0.3$, confirming diffusive dynamics on the sphere. Individual runs show $\alpha$ ranging from $0.55$ to $1.13$, reflecting the different free energy landscapes explored by different realizations. The autocorrelation of $\Theta_{z}$ decays rapidly ($\tau<0.2$), consistent with the white-noise character of Langevin dynamics.
The relaxation dynamics exhibits glassy features: two-stage relaxation (fast dispersion followed by band assembly), metastability, and fluctuation reduction upon structure formation.
Band Orientation Dynamics.
We now analyze the structure normal $\hat{\mathbf{n}}(t)$ as defined in Section 2.5. On the sphere, the band normal undergoes slow diffusive precession on the unit sphere $S^{2}\cong\mathrm{SO}(3)/\mathrm{SO}(2)$, which is precisely the Goldstone manifold for the broken rotational symmetry. Figure 2 presents the complete orientation dynamics.
  • Structure formation: The thickness decreases from $\mathcal{T}\approx 0.62$ (diffuse, nearly isotropic) during formation to $\mathcal{T}\approx 0.29$ (well-defined band) in NESS, confirming dimensional collapse.
  • Isotropic orientation: The mean $\langle n_{z}\rangle\approx 0$ with comparable standard deviations in all components ($\sigma(n_{x})\approx\sigma(n_{y})\approx\sigma(n_{z})\approx 0.1$$0.3$), indicating that the band normal explores all directions on $S^{2}$ without preferred orientation.
  • Long correlation time: The orientation autocorrelation $C(\tau)=\langle\hat{\mathbf{n}}(t)\cdot\hat{\mathbf{n}}(t+\tau)\rangle$ decays to $1/e$ at $\tau_{c}\approx 18$ time units, much longer than the thermal velocity correlation time ($\tau_{v}<0.2$). This separation of timescales confirms the adiabatic nature of orientation dynamics.
  • Cumulative precession: The band normal accumulates approximately $2300^{\circ}$ of angular displacement during the NESS phase ($t=13$ to $50$), corresponding to a mean precession rate of $\sim 61^{\circ}$ per unit time. This continuous exploration of the Goldstone manifold distinguishes our “disorder-induced symmetry breaking” from true spontaneous symmetry breaking where the order parameter would be frozen.
  • Power spectrum: The structure normal power spectrum $S_{n_{z}}(\omega)\propto\omega^{-1.6}$ is shallower than the $1/\omega^{2}$ prediction for type-A Goldstone diffusion modes. This may reflect the bounded nature of $\hat{\mathbf{n}}$ on $S^{2}$.
Figure 2: Sphere $S^{2}$: Band orientation dynamics for $N=400$ particles, $T=0.4$, full time series $t=0$ to $50$. Top row: Band normal components $n_{x}(t)$, $n_{y}(t)$, $n_{z}(t)$ showing slow precession; coral shading indicates formation phase ($t<13$), blue shading indicates NESS. Second row: Band thickness $\mathcal{T}(t)$ decreasing from $\sim 0.6$ to $\sim 0.3$ as the band forms; polar angle $\theta(t)$ of the band normal; cumulative angular displacement showing steady precession at $\sim 61^{\circ}$/time in NESS. Third row: Orientation autocorrelation $C(\tau)=\langle\hat{\mathbf{n}}(t)\cdot\hat{\mathbf{n}}(t+\tau)\rangle$ with correlation time $\tau_{c}\approx 18$; component autocorrelations in NESS; thickness autocorrelation. Bottom row: Power spectra showing $S_{n_{z}}(\omega)\propto\omega^{-1.6}$; 3D trajectory of the band normal on $S^{2}$ demonstrating isotropic exploration of the Goldstone manifold.
Summary: Sphere Dynamics.
On the sphere, dimensional reduction produces a single equatorial band that forms by $t\approx 5$ and persists indefinitely. The cumulative rotation $\Theta_{z}(t)$ exhibits diffusive scaling with MSD exponent $\alpha=0.9\pm 0.3$, consistent with fast, thermally-driven dynamics. The structure normal $\hat{\mathbf{n}}(t)$ (the Goldstone mode variable) precesses slowly and isotropically on $S^{2}$ with correlation time $\tau_{c}\approx 18$, far exceeding the thermal correlation time. Both observables reflect the zero-torque property: no deterministic drift from pairwise forces. The sphere thus provides the cleanest example of disorder-induced symmetry breaking, where the structure forms rapidly but its orientation evolves stochastically on the Goldstone manifold.